At the beginning of the current period, Golden Valley Corporation had balances in Accounts Receivable of $200,700 and in Allowance for Doubtful Accounts of $9,600 (credit). During the period, Golden Valley had credit sales of $802,100 and collections on account of $767,100. Golden Valley uses a perpetual inventory system and determined that the cost of goods sold during the period was $675,100. Golden Valley wrote off as uncollectible accounts receivable of $7,600. However, a $3,300 account previously written off as uncollectible was recovered before the end of the current period. Uncollectible accounts are estimated to total $27,300 at the end of the period. What is the carrying amount of the receivables at the end of the period?
Answer #1

Ending balance of account receivable = 200700+802100-767100-7600 = 228100

Ending balance of allowance for uncollectible = 27300

Carrying amount of receivable = 228100-27300 = 200800